### v1

Abstract
Understanding the nature of primordial fluctuations is critical to our comprehension of the Universe’s early stages. While these fluctuations are known to be nearly scale-invariant, quasi-adiabatic, and nearly Gaussian on large scales, their behavior at smaller scales remains less well-defined and may offer insights into new physics. Recent observations by the NANOGrav, PPTA, EPTA, and CPTA collaborations suggest the presence of a stochastic gravitational wave background, which, while consistent with the contribution from supermassive black hole binaries, also opens the possibility of probing new physics. This paper explores whether this signal could stem from primordial isocurvature and adiabatic fluctuations. We adopt parameterized spectra for both types of fluctuations to fit the observations from the latest NANOGrav data. Furthermore, we employ Bayesian analysis to assess the distinguishability of these models in light of current PTA sensitivities. Our findings indicate that with the capabilities, PTAs cannot conclusively differentiate between isocurvature and adiabatic fluctuations.
keywords:
adiabatic fluctuation, isocurvature fluctuation, pulsar timing array
[…]
04.30.Db, 04.80.Nn, 95.55.Ym
1 Introduction
[…]
Our Universe is characterized by primordial fluctuations that exhibit certain properties on large scales. These fluctuations are remarkably small, nearly scale-invariant, quasi-adiabatic, and almost Gaussian. Their presence is directly observed through the anisotropy pattern of the cosmic microwave background (CMB) [4, 5]. However, our knowledge and constraints regarding the state of the Universe on small scales, where nonlinear physics processes primordial information, are much more limited. When examining large scales, the CMB provides valuable insights, indicating that isocurvature primordial fluctuations likely contribute no more than $1-10\%$ of the total fluctuations [5]. Considering that the measured amplitude of the power spectrum of primordial adiabatic fluctuations is approximately $10^{-9}$, it follows that the power spectrum of isocurvature fluctuations may have an amplitude of less than $10^{-10}$ on large scales. However, the situation changes when we shift our focus to small scales, where the constraints imposed by the CMB no longer hold. On these smaller scales, our most promising avenue for constraining the amplitude and nature of primordial fluctuations lies in the study of primordial black holes (PBHs) [6, 7, 8] and induced gravitational waves (IGWs) [9, 10, 11, 12, 13, 14, 15]. To investigate the early Universe’s nature further, pulsar timing arrays (PTAs) prove to be indispensable tools. PTAs possess sensitivity in the nHz frequency band, enabling them to probe the characteristics of the early Universe in a unique and valuable manner.
Recently, several PTA collaborations, including the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [16, 17], the Parkers PTA (PPTA) [18, 19], the European PTA (EPTA) in conjunction with the Indian PTA (InPTA) [20, 21], and the Chinese PTA (CPTA) [22], have collectively unveiled compelling evidence supporting the existence of a stochastic gravitational wave background (SGWB) within the frequency range of approximately $1-10\,\mathrm{nHz}$, which attracts lots of implications [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61]. While the inferred amplitude and spectrum of the GW background align with astrophysical expectations for a signal originating from the population of supermassive black hole binaries (SMBHBs), the possibility of new physics within this observational window remains a tantalizing prospect. One plausible explanation for the observed SGWB is that it arises from IGWs stemming from primordial fluctuations [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99]. Furthermore, cosmological phenomena such as first-order phase transitions [100, 101, 102, 103, 104, 105, 106, 107, 108] and the presence of topological defects like cosmic strings [109, 110, 111, 112, 113, 114, 115, 116] and domain walls [117, 118, 119] could also serve as sources for generating the SGWB. The SGWB generated by these cosmological sources holds immense potential for advancing our understanding of physics beyond the Standard Model and providing invaluable insights into the early stages of the Universe.
In this study, our goal is to investigate the potential explanations for the observed signals through IGWs stemming from both primordial isocurvature and adiabatic fluctuations. To achieve this, we utilize parameterized primordial spectra for both types of fluctuations to effectively match the latest observations reported by the NANOGrav. Moreover, we calculate the Bayes factors to compare different models. Our analysis indicates that, with the current sensitivity of PTAs, it is challenging to differentiate between the contributions of isocurvature and adiabatic fluctuations. The structure of this paper is outlined as follows: In Section 2, we introduce the general formalisms for IGWs arising from both types of primordial fluctuations. Section 3 details the methodology employed in our data analysis and presents the findings using the latest NANOGrav data set, along with the computed Bayes factors for comparing different models. We conclude by summarizing our results and discussing their implications in Section 4.
2 IGWs from primordial isocurvature fluctuations and primordial adiabatic fluctuations
For the SGWB to be detectable, it is crucial for the primordial perturbations to undergo significant amplification compared to the fluctuations observed in CMB experiments. The amplitude and shape of the SGWB spectrum are influenced by various factors associated with the primordial fluctuations. These factors encompass the power spectrum and non-Gaussianities linked to the fluctuations, as well as the initial conditions of the primordial perturbations. Together, these elements shape the characteristics of the SGWB spectrum, determining both its amplitude and shape. In this section, we will provide an overview of the general formalisms concerning IGWs arising from primordial isocurvature fluctuations and primordial adiabatic fluctuations. The formalism for analytically calculating GWs induced by primordial adiabatic fluctuations during the radiation-dominated epoch was initially established in Refs. [120, 121, 9, 10, 122, 14] and reviewed in Ref. [123]. Conversely, the formalism for analytically calculating GWs induced by primordial isocurvature fluctuations during the radiation-dominated epoch was introduced in Refs. [124] and reviewed in Ref. [125].
The characterization of SGWBs today often involves describing their energy density per logarithmic frequency interval relative to the critical density $\rho_{c}$,
$\Omega_{\mathrm{GW}}(k)\equiv\frac{1}{\rho_{c}}\frac{\mathrm{d}\rho_{\mathrm{GW}}(k)}{\mathrm{d}\ln k}.$ (1)
After generation, the energy density of GWs evolves in the same way as radiation. Utilizing this property, it is straightforward to determine the energy density of GWs at present. Taking the late-time limit during the radiation-dominated era, the relation of the GW energy density at present and the generation is given by [126]
$\Omega_{\rm GW}(k)$ $=\Omega_{r,0}\left(\frac{g_{*,r}(T)}{g_{*,r}(T_{0})}\right)\left(\frac{g_{*,s}(T)}{g_{*,s}(T_{0})}\right)^{-4/3}\Omega_{\rm GW,c}(k)\,$
$\approx 0.4\left(\frac{g_{*,r}(T)}{80}\right)\left(\frac{g_{*,s}(T)}{80}\right)^{-4/3}\Omega_{r,0}\,\Omega_{\rm GW,c}(k)\,$ (2)
where $\Omega_{r,0}$ is the density fraction of radiation in the present epoch, $g_{*,r}(T_{c})$ and $g_{*s}(T_{c})$ are the effective number of degrees of freedom in the energy density and entropy, respectively. The values today are denoted as $g_{*,r}(T_{0})=3.36$ and $g_{*,s}(T_{0})=3.91$. For the temperature dependence of $g_{*,r}(T)$ and $g_{*,s}(T)$, see Ref. [127].
Following the Refs. [14, 122], the energy density of IGWs arising from primordial adiabatic fluctuations at the epoch of matter-radiation equality can be presented as
$\Omega_{\mathrm{GW,c}}(k)=\int_{0}^{\infty}\mathrm{d}v\int_{|1-v|}^{1+v}\mathrm{d}u\mathcal{T}_{\zeta}(u,v){\mathcal{P}}_{\zeta}(ku){\mathcal{P}}_{\zeta}(kv),$ (3)
where ${\mathcal{P}}_{\zeta}$ is the primordial power spectrum of curvature perturbations and
$\mathcal{T}_{\zeta}(u,v)=$ $\frac{3}{1024v^{8}u^{8}}\left[4v^{2}-\left(v^{2}-u^{2}+1\right)^{2}\right]^{2}\left(v^{2}+u^{2}-3\right)^{2}$ (4)
×$\times\bigg{\{}\left[\left(v^{2}+u^{2}-3\right)\ln\left(\left|\frac{3-(v+u)^{2}}{3-(v-u)^{2}}\right|\right)-4vu\right]^{2}$
$+\pi^{2}\left(v^{2}+u^{2}-3\right)^{2}\Theta(v+u-\sqrt{3})\bigg{\}}.$
Under the assumption of Gaussian isocurvature fluctuations, we can derive a concise expression for the spectral density of IGWs originating from primordial isocurvature fluctuations [124]. This expression is given by
[…]
Here, we define $\kappa\equiv k/k_{\rm eq}$, where $k_{\rm eq}$ represents the comoving wavenumber at radiation-matter equality. ${\mathcal{P}}_{s}(k)$ corresponds to the dimensionless spectrum of isocurvature fluctuations. The transfer function $\mathcal{T}_{s}$ is given by
$\mathcal{T}_{s}(\kappa,u,v)=\frac{1}{3}\left(\frac{4v^{2}-(1-u^{2}+v^{2})^{2}}{4uv}\right)^{2}\left(I_{c,\infty}^{2}(\kappa,u,v)+I_{s,\infty}^{2}(\kappa,u,v)\right)\,,$ (6)
where
$I_{c,\infty}(\kappa,u,v)$ $=\frac{9}{32u^{4}v^{4}\kappa^{2}}\Bigg{\{}\left(-3+v^{2}\right)\left(-3+v^{2}+2u^{2}\right)\ln\left|1-\frac{v^{2}}{{3}}\right|$
$-3u^{2}v^{2}+\left(-3+u^{2}\right)\left(-3+u^{2}+2v^{2}\right)\ln\left|1-\frac{u^{2}}{{3}}\right|$
$-\frac{1}{2}\left(-3+v^{2}+u^{2}\right)^{2}\ln\left[\left|1-\frac{(u+v)^{2}}{{3}}\right|\left|1-\frac{(u-v)^{2}}{{3}}\right|\right]\Bigg{\}}\,,$ (7)
and
$I_{s,\infty}(\kappa,u,v)$ $=\frac{9\pi}{32u^{4}v^{4}\kappa^{2}}\Bigg{\{}9-6v^{2}-6u^{2}+2u^{2}v^{2}$
$+\left(3-u^{2}\right)\left(-3+u^{2}+2v^{2}\right)\Theta\left(1-\frac{u}{\sqrt{3}}\right)$
$+\left(3-v^{2}\right)\left(-3+v^{2}+2u^{2}\right)\Theta\left(1-\frac{v}{\sqrt{3}}\right)$
$+\frac{1}{2}\left(-3+v^{2}+u^{2}\right)^{2}\left[\Theta\left(1-\frac{u+v}{\sqrt{3}}\right)+\Theta\left(1+\frac{u-v}{\sqrt{3}}\right)\right]\Bigg{\}}\,.$ (8)
To maintain a high level of model independence, we adopt an approach that avoids selecting a specific inflation model capable of generating an enhanced spectrum $\mathcal{P}_{\zeta/s}$. Instead, we focus on three characteristic templates for $\mathcal{P}_{\zeta/s}$ that encompass a range of possibilities typically observed in realistic models. By disregarding the microphysics of inflation, we aim to provide a broader analysis that accommodates various scenarios. The following templates are considered:
  • Adiabatic/isocurvature delta:
    $\mathcal{P}_{\zeta/s}\left(k\right)=A\,\delta\left(\ln k-\ln k_{*}\right)\,.$ (9)
  • Adiabatic/isocurvature lognormal:
    $\mathcal{P}_{\zeta/s}\left(k\right)=\frac{A}{\sqrt{2\pi}\,\Delta}\,\exp\left[-\frac{1}{2}\left(\frac{\ln k-\ln k_{*}}{\Delta}\right)^{2}\right]\,.$ (10)
  • Adiabatic/isocurvature box:
    $\mathcal{P}_{\zeta/s}\left(k\right)=A\,\Theta\left(\ln k_{\rm max}-\ln k\right)\Theta\left(\ln k-\ln k_{\rm min}\right)\,.$ (11)
3 Data analyses and results
Table 1: Parameters along with their respective prior distributions utilized in the Bayesian inference analysis, where $\mathcal{U}$ denotes a uniform distribution. Results are reported as the median and the $90\%$ credible interval, with symmetrical tails.
Model Parameter Description Prior Results
$M_{1}$ adiabatic delta
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(-5,5)$ $0.56^{+1.83}_{-1.99}$
$\log_{10}(f_{*}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-15,-2)$ $-4.28^{+2.02}_{-2.45}$
$M_{2}$ adiabatic lognormal
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(-2,2)$ $0.21^{+1.47}_{-1.52}$
$\log_{10}(f_{*}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-8,-2)$ $-4.97^{+2.58}_{-2.31}$
$\Delta$ Width of the power spectrum. ${\mathcal{U}}(0.01,5)$ $2.35^{+1.70}_{-1.98}$
$M_{3}$ adiabatic box
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(-3,3)$ $-1.26^{+2.28}_{-0.61}$
$\log_{10}(f_{\mathrm{min}}/\mathrm{Hz})$ Minimum frequency of the power spectrum. ${\mathcal{U}}(-9,-4)$ $-7.37^{+1.86}_{0.04}$
$\log_{10}(f_{\mathrm{max}}/\mathrm{Hz})$ Maximum frequency of the power spectrum. ${\mathcal{U}}(-9,-2)$ $-4.60^{+2.34}_{-3.45}$
$M_{4}$ isocurvature delta
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(1,50)$ $25.3^{+5.9}_{-6.9}$
$\log_{10}(f_{*}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-15,-2)$ $-4.25^{+2.03}_{-2.44}$
$M_{5}$ isocurvature lognormal
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(10,50)$ $21.7^{+5.2}_{-4.3}$
$\log_{10}(f_{*}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-8,-2)$ $-4.95^{+2.64}_{-2.18}$
$\Delta$ Width of the power spectrum. ${\mathcal{U}}(0.01,5)$ $1.16^{+0.79}_{-1.11}$
$M_{6}$ isocurvature box
$\log_{10}A$ Amplitude of the power spectrum. ${\mathcal{U}}(10,30)$ $18.3^{+4.9}_{-1.3}$
$\log_{10}(f_{\mathrm{min}}/\mathrm{Hz})$ Pivot frequency. ${\mathcal{U}}(-9,-4)$ $-7.10^{+1.54}_{-0.45}$
$\log_{10}(f_{\mathrm{max}}/\mathrm{Hz})$ Width of the power spectrum. ${\mathcal{U}}(-9,-2)$ $-4.41^{+2.17}_{-2.44}$
In this study, we analyze the NANOGrav 15-year data set [17] to infer the parameters of our model. Our approach particularly leverages the amplitudes of the free spectrum as determined by NANOGrav while accounting for the spatial correlations consistent with the Hellings-Downs [128] curve. The observational sensitivity of a PTA begins at a frequency equivalent to the inverse of the total observation time, $1/T_{\mathrm{obs}}$, where $T_{\mathrm{obs}}=16.03\,\mathrm{yr}$ represents the duration covered by the NANOGrav 15-year data set. We incorporate the posterior distributions of the free spectrum for the $14$ frequency components reported by NANOGrav [16] in their analysis of the SGWB signal.
Our analysis begins by examining the posterior time delay data, $d(f)$, as provided by NANOGrav. The relationship between this time delay and the power spectrum, $S(f)$, is
$S(f)=d(f)^{2}\,T_{\mathrm{obs}}.$ (12)
Utilizing the time delay data allows us to determine the SGWB energy density spectrum by
$\hat{\Omega}_{\mathrm{GW}}(f)=\frac{2\pi^{2}}{3H_{0}^{2}}f^{2}h_{c}^{2}(f)=\frac{8\pi^{4}}{H_{0}^{2}}T_{\mathrm{obs}}f^{5}d^{2}(f),$ (13)
in which $H_{0}$ represents the Hubble constant and $h_{c}(f)$ is the characteristic strain, calculated as
$h_{c}^{2}(f)=12\pi^{2}f^{3}S(f).$ (14)
At each observed frequency $f_{i}$, the kernel density estimate, $\mathcal{L}_{i}$, is derived from the associated $\hat{\Omega}_{\mathrm{GW}}(f_{i})$ posteriors. Consequently, the total log-likelihood is the sum across all frequencies of the individual log-likelihoods [68, 37, 85, 79]:
$\ln\mathcal{L}(\Lambda)=\sum_{i=1}^{14}\ln\mathcal{L}_{i}(\Omega_{\mathrm{GW}}(f_{i},\Lambda)),$ (15)
where $\Lambda$ is a collection of model parameters that will be inferred from the PTA data.
For the parameter space exploration, we utilize the dynesty sampler [129] incorporated within the Bilby package [130, 131]. The conversion between the wavenumber $k$ and the frequency $f$ is expressed by
$f=\frac{k}{2\pi}\simeq 1.6\,\mathrm{nHz}\left(\frac{k}{10^{6}\,\mathrm{Mpc}^{-1}}\right).$ (16)
which provides the necessary link for interpreting spectral features in terms of their spatial scales. In this work, we consider six models: the adiabatic delta model ($M_{1}$), adiabatic lognormal model ($M_{2}$), adiabatic box model ($M_{3}$), isocurvature delta model ($M_{4}$), isocurvature lognormal model ($M_{5}$), and isocurvature box model ($M_{6}$). The selected priors for the model parameters are comprehensively listed in Table 1.
Figure 1: Marginalized one- and two-dimensional posterior distributions for the model parameters of the adiabatic delta (upper panel) and isocurvature delta (lower panel) models, derived from the NANOGrav 15-year data set. The contours in the 2D distributions represent the $1\sigma$, $2\sigma$, and $3\sigma$ credible intervals, respectively.
The posterior distributions for the parameters of the models $M_{1}$ through $M_{6}$ are illustrated in Figs. 1-3. Our analysis indicates that a successful explanation of the NANOGrav signal with IGWs requires a substantial amplitude $A$ for the primordial scalar power spectrum. This inference is supported by computing the lower limits of the $95\%$ Bayesian credible intervals for the amplitude $A$, derived from the one-dimensional marginalized posteriors over the regions of the highest posterior density. For models $M_{1}$ through $M_{6}$, we find the following lower limits: $\log_{10}A\gtrsim-1.43$, $\log_{10}A\gtrsim-1.31$, $\log_{10}A\gtrsim-1.88$, $\log_{10}A\gtrsim 18.3$, $\log_{10}A\gtrsim 17.4$, and $\log_{10}A\gtrsim 17.0$, respectively. Additionally, the amplification in scalar perturbations must be confined to specific scales to ensure that the resulting IGW signal is within the sensitivity range of PTAs. This constraint imposes specific bounds on the characteristic frequencies $f_{\mathrm{min}}$, $f_{\mathrm{max}}$, and $f_{*}$, which are evident from Figs. 1-3 and are concisely presented in Table 1.
Figure 2: Same as Fig. 1, but for the adiabatic lognormal (upper panel) and isocurvature lognormal (lower panel) models.
A significant observation is that the posterior distributions for characteristic frequencies, $f_{\mathrm{min}}$, $f_{\mathrm{max}}$, and $f_{*}$, extend to values well above the upper limit of the NANOGrav sensitivity range. This implies that the NANOGrav data align more closely with the low-frequency end of the IGW spectrum. The IGW spectrum continues to rise beyond the NANOGrav frequencies, reaching its peak at $f\gg 1\mathrm{nHz}$.
This behavior is further corroborated by the flat trends observed in the 2D posterior distributions of the amplitude $A$ with both $f_{\mathrm{min}}$ and $f_{*}$, as illustrated in Figs. 1-3. These flat directions in the posterior space indicate that increasing $A$ simultaneously with $f_{\mathrm{min}}$ or $f_{*}$ shifts the GW spectrum’s peak to higher frequencies. Crucially, this shift does not alter the low-frequency tail within the PTA band, thereby preserving the fit’s consistency with the NANOGrav data.
Table 2: Evaluation of model selection scores quantified by the Bayes factor as described in Ref. [132].
   BF Strength of evidence
$<1$ Negative
$1-3$ Not worth more than a bare mention
$3-20$ Positive
$20-150$ Strong
$>150$ Very strong
Upon examination, the 2D posterior distributions for the amplitude ($A$) and characteristic frequency ($f_{*}$) under both delta and lognormal power spectra exhibit a remarkable resemblance. However, it is noted that the distribution corresponding to the lognormal model is marginally wider. This observation aligns with the theoretical understanding that delta and lognormal models are hierarchically related; specifically, the delta power spectrum is a particular instance of the lognormal spectrum as the width parameter $\Delta$ approaches zero. Consequently, the somewhat expanded posterior distribution for $A$ and $f_{*}$ illustrated in Fig. 2 mirrors the additional parameter space afforded by the lognormal model, which in turn introduces an incremental degree of parametric freedom. The broader scope of the lognormal model’s distribution aptly captures this extra complexity.
To assess which model is most strongly supported by the observational data, we compute the Bayes factor for the competing models. The Bayes factor offers a statistical metric for comparing two mutually exclusive models, $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$. It measures the strength of evidence in favor of one model over the other and is defined as
$\mathrm{BF}_{ij}\equiv\frac{\mathcal{Z}_{i}}{\mathcal{Z}_{j}},$ (17)
where $\mathcal{Z}_{i}$ and $\mathcal{Z}_{j}$ represent the evidence for models $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$, respectively. The evidence, denoted by $\mathcal{Z}$, is calculated by integrating the likelihood $\mathcal{L}(\textbf{d}|\Lambda)$ of the data d given the parameters $\Lambda$, weighted by the prior probability distribution $\pi(\Lambda)$ of the parameters, namely
$\mathcal{Z}\equiv\int\mathrm{d}\Lambda\,\mathcal{L}(\textbf{d}|\Lambda)\,\pi(\Lambda).$ (18)
Table 2 summarizes the Bayes factor interpretations for model comparison [132].
Figure 3: Same as Fig. 1, but for the adiabatic box (upper panel) and isocurvature box (lower panel) models.
The calculation of Bayes factors among six competing models yields the following results:
$\mathrm{BF}_{ij}=\left(\begin{array}[]{ccccccc}&M_{1}&M_{2}&M_{3}&M_{4}&M_{5}&M_{6}\\ M_{1}&1.00&0.54&0.87&1.78&8.50&2.76\\ M_{2}&1.86&1.00&1.62&3.32&15.85&5.15\\ M_{3}&1.18&0.62&1.00&2.04&9.76&3.17\\ M_{4}&0.56&0.30&0.49&1.00&4.77&1.55\\ M_{5}&0.12&0.06&0.10&0.21&1.00&0.32\\ M_{6}&0.36&0.19&0.32&0.64&3.07&1.00\end{array}\right).$ (19)
Upon analysis, model $M_{2}$ exhibits the highest Bayes factor, in contrast to model $M_{5}$, which has the lowest. Remarkably, the Bayes factor comparing $M_{2}$ to $M_{5}$ is $\mathrm{BF}_{25}=15.85$, indicating positive evidence rather than significant evidence in favor of $M_{2}$ over $M_{5}$. Therefore, the interpretation of the NANOGrav 15-year dataset concerning the common-spectrum process and the associated Hellings-Downs correlation remains nuanced. The Bayes factors, when considered collectively, provide evidence that is not strong enough to decisively favor any single model over the others. Consequently, selecting the optimal model among $M_{1}$ through $M_{6}$ is challenging, given the narrow range of their Bayes factors and the level of evidence they present.
4 Summary and discussion
In this study, we have investigated the possibility of explaining the SGWB observed by PTAs through IGWs arising from primordial isocurvature and adiabatic fluctuations. By analyzing the latest data from the NANOGrav collaboration, we have computed the Bayes factors comparing various models. Our findings indicate that the current sensitivity of PTAs is insufficient to distinguish between isocurvature and adiabatic fluctuations. The analysis of the NANOGrav data, while providing valuable insights into the nature of the SGWB, does not allow us to conclusively determine the origin of the observed PTA signal. This result highlights the challenges involved in unraveling the primordial origins of the SGWB and the need for further advancements in sensitivity and data analysis techniques.
While our current analysis suggests that PTAs cannot yet distinguish between isocurvature and adiabatic fluctuations, we anticipate that future advancements in observational capabilities and data analysis techniques will gradually improve our ability to probe the primordial origins of the SGWB. The continued progress in constructing next-generation radio telescopes, such as the Square Kilometre Array [133], promise to provide unprecedented sensitivity and frequency coverage, enabling more precise measurements of the SGWB and its properties.
Furthermore, ongoing efforts to refine theoretical models of the early Universe and the generation of primordial fluctuations will contribute to our understanding of the observed PTA signal. Refining our understanding of the nonlinear physics involved in the evolution of small-scale fluctuations will be crucial for accurately modeling the IGWs and distinguishing them from other sources contributing to the SGWB.
In conclusion, our study emphasizes the importance of the SGWB as a powerful probe of the early Universe and highlights the potential of PTAs in unraveling the primordial origins of GWs. While the current sensitivity of PTAs is insufficient to distinguish between isocurvature and adiabatic fluctuations, future advancements in observational capabilities and theoretical modeling, combined with multi-PTA and multi-messenger approaches, hold the promise of shedding light on the nature and origin of the observed PTA signal.
[…]
ZCC is supported by the National Natural Science Foundation of China (Grant Nos. 12247176 and 12247112) and the China Postdoctoral Science Foundation Fellowship (Grant No. 2022M710429). LL is supported by the National Natural Science Foundation of China (Grant Nos. 12247112 and 12247176) and the China Postdoctoral Science Foundation Fellowship (Grant No. 2023M730300).
[…]
The authors declare that they have no conflict of interest.